Systems Staff Meeting Notes 10/30/00

Systems Staff Meeting Notes
October 30, 2000

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Debra Harris, Winston Harris, Martha Hruska, Bobbie Meng, Suzy Shaw, Hilton Soto, Geoffrey Stewart, Ying Tang
Announcements
  1. Bring treats tomorrow for the Liaison Halloween Party.
  2. Suzy was locked out of her office this morning - the only person that could gain access was Maybell. Suspected to be more of the "security" implementation.
  3. UF won this weekend.
  4. We didn't win the lottery - it goes up to $35 million on Wednesday.
Review Minutes of 10/23/00
  1. RIR #7 - Wireless - One laptop is marked for Resource Services, but the other is to go to Special Collections. Will distribute to Resource Services today.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Order for 20 more Dells was taken to Barbara this morning - hopefully with the correct specs! The order includes the specific instructions to deliver before 3:00 PM, inside delivery. If they show after this time, we are to refuse delivery. They are also supposed to call the day before delivery. Debra will send an email to Facilities, making them aware of the delivery restrictions.
  2. Bar Code Readers - no complaints so far, will probably need to order more. Bobbie notes that we do have some long-range scanners in the vault - maybe we can return them to the company & have them swapped out.
  3. Laptops - only one wireless for Resource Services. Nine with Windows 2000 are for liaison training; two that will be used for checkout when swapped with the Toshibas will be downgraded to Win'98. The Toshibas need to have the drives formatted, then surveyed.
  4. Have the bags for the laptops been ordered yet? Unsure, will check on it.
  5. Andy put in the order for the part for the NT8 server.
  6. Erich - super scanner - Geoffrey checked on the specs & emailed them to Erich. Estimated at half of what was originally planned. Erich & Will will go to Orlando for a demo of the 4500 as soon as Erich can set it up.
  7. Server Room proposal - can scale back to two servers for right now: cluster for Exchange, tape rack, power supplies and one tape drive.
Liaisons
  1. Training session this morning went well. 10 of the 16 scheduled to show made it to the session. The handouts were clear; session started at 8:30 AM and was concluded by 10:30 AM.
  2. No evaluation sheet was handed out. Meeting scheduled for this Thursday, will ask for feedback.
  3. Next session scheduled for Monday, Nov. 6 - from 1-3PM. Geoffrey will get with Debra about updating the Liaison section of the Systems Web page.
  4. May schedule these training sessions once a semester to catch the new liaisons.
Networking
  1. Net Agent Server - Will needs to generate the specs. The plan as of last week has changed - trouble with moving Exchange voided the plan. Will order a server similar to NT2 for Net Agent. Suzy goes to Net Agent training in Atlanta next week. Training for others will be some time in December, so we'll need the server working & effective by the start of classes in January.
  2. Print server - need to move files on NT2 to NT7. NT7 magically powered back on this weekend. Will can spec it out and have another server ordered today. Should be similar to NT2. We do have some older servers in the vault that could be brought up.
  3. NT 5 died sometime Friday. It is now active again, however, it is at Service Pack 4.
  4. Public Printing - Xerox - meeting last week, discussed possible dates to install the NT version - maybe Christmas break. They need to install patch to allow for Cows1 login. They are waiting on parts for the server. Music will be the first department to test it - also the first section of West. May try out the system the week of November 9th - the week prior to Homecoming. If it doesn't work, will try again at Christmas. Will keep both NT & Novell up while testing.
  5. Proxy Server - at least three outages have been reported. John Ashcraft called in the problem. Think it may be a traffic issue. Should probably start a log by the machine to document the time of day that outages are occurring. When the server freezes, the quickest fix is to hit control-alt-delete.
  6. Cataloger's Desktop - Will has it on his to-do list, he is looking for his notes. If he is unable to locate them, he has a name and number to call for assistance.
  7. Rewiring Project - supposed to start later this week. Documents is finished. Bobbie asks if the extra pod in room 300 will have phones. Unknown.
  8. Pull from room 209-210 to room 310 is complete.
  9. 1 pull in MSL forthcoming.
  10. 200 East - haven't moved the fiber yet, but need a rack.
  11. Bobbie foresees electrical problems with the 2nd floor East Com Closet once the phones go in there. Brad is aware of the situation.
  12. Will pull new circuit for the 1st floor Com closet.
  13. Encumbrance for MSL - work that was already done - may move funds over for additional pulls. Helen Jane's office is where the pull should go.
  14. Switches - can't order more, need to rearrange & reconfigure the ones we have.
  15. What about rewiring rooms 100 & 148 in West? It was not on the FCLA funding wish list. Both rooms need to be done - they are on Hubs right now.
  16. Wireless - only the third floor is wired now. It is useable on the 3rd & parts of the 2nd & 4th floors. The laptop delivered to Resource Services can be both wireless and standard.
Software Projects Status
  1. Film Log, DLC Registry & FULoader are all pending outside forces.
  2. Will deploy the read-only IR database to Facilities this week for testing.
  3. Facilities Supply Inventory database - is currently in Access. The HD it was on died. They may have one backup file of it. We may eventually need to put the database on the server.
  4. Hilton has been working on mods to the IRDB. He has finished two of three; the 3rd change is harder and will require a short down time. Should be complete in two weeks.
  5. Accounting requested a new report & Special Collections needed a change on a dynamic Web page taken care of last week.
DP Coordinator
  1. Tape Library List - Dolores wanted a printout other than the text file that Winston compiled. Debra gave Will an old 1997 computer printout of the volser listing. He will pass it on to Dolores. Martha will speak to John Ingram about proceeding with this project.
Calendar
  1. Suzy out November 6-10 (Net Agent Training in Atlanta).
  2. Geoffrey out November 23.
  3. Will & Hilton both out Nov. 20-22 & 27.


Back to the Systems Homepage

Last updated November 6, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu